### v1

# Two-Electron Correlations in the Metallic Electron Gas

Abstract
We present high-precision ab initio calculations of the four-point vertex function for the three-dimensional uniform electron gas using variational diagrammatic Monte Carlo. From these results, we extract Landau parameters that demonstrate a density-driven crossover from underscreening to overscreening. Guided by our numerical data, we propose a charge-based Kukkonen–Overhauser effective interaction within the local-density approximation, supplemented by a small s-wave correction (sKO+), which accurately captures the electron–electron scattering amplitude. Using our numerically determined scattering amplitude, together with the sKO+ ansatz, we compute the electron-electron contribution to the thermal resistivity, demonstrating excellent agreement with experimental measurements in simple metals.
thanks: These two authors contributed equally to this paper.thanks: These two authors contributed equally to this paper.
Introduction.— The collective behavior of interacting electrons gives rise to some of the most profound and challenging problems in condensed matter physics [1], including high-temperature superconductivity [2], strange metallicity [3], and the emergence of topological phases [4]. The uniform electron gas (UEG)—an idealized model of electrons moving in a uniform neutralizing background—provides the quintessential canvas for understanding correlations originating solely from the Coulomb interaction [5]. While the single-particle properties of this system have been extensively studied [6], a comprehensive understanding of two-electron correlations has remained elusive despite their fundamental importance.
The key to this problem lies in the four-point vertex function, which is central to understanding Fermi liquid behavior [7, 8], pairing instabilities [9], and electron-electron scattering rates that govern transport phenomena [10]. A first-principles determination has historically been impeded by its complex dependence on multiple momentum and frequency variables. The immense computational cost associated with this high dimensionality poses a severe challenge, forcing many theoretical frameworks to rely on approximations, such as low-order truncations in diagrammatic methods or the local approximations in dynamical mean-field extensions and post-GW schemes [11, 12, 13, 14]. At the same time, conventional quantum Monte Carlo techniques also face challenges, as the vertex function captures subtle many-body correlations that are difficult to extract directly from the static ground-state densities typically sampled in these simulations. This theoretical gap is particularly pressing today, as advances in two-electron photoemission and multidimensional electronic spectroscopies begin to access many-electron correlations experimentally [15, 16, 17].
Figure 1: The electron-electron contribution to the thermal resistivity $W_{ee}$ as a function of density parameter $r_{s}$ in the metallic regime derived through our numerical results (blue circles) through the variational diagrammatic Monte Carlo (VDMC) method and some theoretical models as the random phase approximation (RPA), s-p approximation [18, 19], the original Kukkonen-Overhauser interaction [20] (KO±) and the one with charge channel only (KO+). For comparision, we plot the experimental data (black circle) for some simple metals [21, 22, 23, 24]. The solid red line presents our proposed sKO+ model, the KO+ model with s-wave corrections, which shows good agreement with the VDMC results and the experimental measurements. The inset shows the results of the electron-electron scattering rate.
The consequences of this theoretical gap are particularly stark in the study of transport phenomena, presenting a long-standing challenge to the first-principles description of thermal conductivity in simple metals. Here, e-e scattering is understood to be a primary contributor to the experimentally observed deviations from the Wiedemann–Franz law [25, 26, 21, 22, 23, 24], an effect that conventional density functional theory (DFT) struggles to account for [27]. Previous attempts to resolve this issue using many-body theories, such as the random phase approximation (RPA)  [28, 29] and the Kukkonen-Overhauser (KO) interaction [29, 20], have also resulted in thermal resistivity values that disagree with experimental measurements. Consequently, finding a theoretical ansatz for the effective electron-electron interaction that yields precise quantitative agreement with experiment remains an open challenge.
In this work, we address this challenge using the variational diagrammatic Monte Carlo (VDMC) method [30, 31, 32, 33], which combines a variational principle with high-order, controlled stochastic summation of Feynman diagrams [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. For the 3D UEG, we compute the four-point vertex up to sixth order, comprehensively characterizing two-electron correlations and extracting the Landau parameters $F_{0,1,2}^{s,a}$. This provides a long-sought, first-principles quantitative determination of these key Fermi-liquid parameters. Our results reveal a crossover from underscreening to overscreening [45, 46, 47], where collective screening drives the static dielectric function negative, effectively inverting the test-charge interaction from repulsive to attractive. This transition is characterized by the Landau parameter $F_{0}^{s}$ approaching 1-1, signaling a divergence in electronic compressibility and the onset of a strongly correlated regime.
Furthermore, we calculate the full e-e scattering amplitude with high precision across the entire Fermi surface. Crucially, these results serve as the foundation for our proposal of a robust effective interaction: a charge-based KO model within the local density approximation (LDA), supplemented by an s-wave correction (sKO+). This ansatz effectively resolves the discrepancies found in widely used approximations, such as the RPA and the KO interaction. Notably, our sKO+ ansatz offers a consistent description of electron correlations across all energy scales, extending the high precision rigorously established on the Fermi surface.
Crucially, our first-principles results for the e-e scattering amplitude enable a direct calculation of the thermal resistivity, resolving the long-standing discrepancy between previous theoretical approximations and experimental measurements. As shown in Fig. 1, the resistivity derived from our calculated scattering amplitude is in good agreement with experimental data [21, 22, 23, 24]. Moreover, the thermal resistivity calculated using our proposed sKO+ model also closely matches the experimental values, confirming the validity of our comprehensive approach.
Methods.— We employ the VDMC framework  [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40], which combines a robust variational principle with stochastic sampling of Feynman diagrams. This approach enables controlled evaluation of high-order perturbative contributions to the four-point vertex function of the uniform electron gas.
The central idea is to reorganize many-body perturbation theory around a variationally optimized screened interaction. Counterterms are introduced to preserve the exact physics of the original Hamiltonian, ensuring that the expansion remains unbiased. The diagrammatic series is then sampled as a computational graph, where counterterm contributions are obtained through Taylor-mode automatic differentiation.
Our implementation incorporates automated high-order renormalization and GPU-accelerated integration routines [41], allowing efficient evaluation of high-dimensional diagrams up to sixth order. This combination provides unprecedented computational reach for effective field theories in strongly interacting systems.
A key feature of VDMC is the presence of systematic cancellations between competing diagrams, which accelerates convergence of physical observables. For quantities such as the Landau parameters, this leads to remarkably stable results: for instance, at $r_{s}=4$ the combined parameters converge by sixth order, with uncertainties estimated from the spread of the last three expansion orders which is shown in Fig. S1 in the supplemental material.
Figure 2: Numerical results of the symmetric (left panel) and antisymmetric (right panel) Landau parameters versus $r_{s}$, which confirms the rapid convergence of the Legendre expansion. The linear decrease of $F_{0}^{s}$ with $r_{s}$, which approaches 1-1 at $r_{s}^{c}\approx 5.7$, reveals the crossover of the Coulomb screening effect from underscreening to overscreening [45]. For comparison, the orange dash-dot line shows the $F_{0}$ calculated from the RPA vertex. The violet dash-dot line presents the dimensionless static exchange-correlation (XC) kernel $f_{\rm XC}(q=0)N_{F}$ for the uniform electron gas derived from the PW92 energy functional [48], which demonstrates the approximate equality of the XC kernel and the zeroth Landau parameter. The corresponding numerical data are listed in Table S1 of the Supplemental Material.
Results.— The 3D UEG is modeled by the reduced temperature $\Theta=T/T_{F}$ and the density parameter, the Wigner-Seitz radii $r_{s}=(\frac{3}{4\pi n})^{1/3}$ with Coulomb interaction $v_{q}=4\pi e^{2}/q^{2}$, where $T_{F}$ is the Fermi temperature, $n$ is the electron density. We investigate the parameter regimes defined by $\Theta=0.025$ and $0.05$, with $r_{s}$ ranging from 1 to 6. Within Landau-Silin theory [49, 50], the effective short-range interaction between quasiparticles on the Fermi surface $f_{\sigma\sigma^{\prime}}(\theta)$, which encapsulates many-body correlations, is obtained from the small-frequency and long-wavelength limit of the four-point vertex diagram $\Gamma_{4}$ as
$f_{\sigma\sigma^{\prime}}(\theta)=z^{2}\lim_{\mathbf{q}\to 0}[\Gamma_{4}^{\sigma\sigma^{\prime}}(\mathbf{k_{1}}\omega_{-1},\mathbf{k}_{2}\omega_{0};\mathbf{k_{1}+q}\omega_{0})-v_{q}].$ (1)
Here, $\theta$ is the angle between the incoming momentum $\mathbf{k}_{1}$ and $\mathbf{k}_{2}$ on the Fermi surface, $\omega_{0}=i\pi T,\omega_{-1}=-i\pi T$ are the Fermionic Matsubara frequency and $z$ is the quasiparticle renormalization factor derived from the calculation of self-energy [41]. Meanwhile, the two-electron scattering amplitude, $a_{\sigma\sigma^{\prime}}(\theta,\phi)$, is related to the four-point vertex for elastic scattering of quasiparticles on the Fermi shell as:
$a_{\sigma\sigma^{\prime}}(\theta,\phi)=z^{2}\Gamma_{4}^{\sigma\sigma^{\prime}}(\mathbf{k}_{1}\omega_{0},\mathbf{k}_{2}\omega_{0};\mathbf{k}_{3}\omega_{0}),$ (2)
where $\phi$ is the angle between the plane defined by $(\mathbf{k}_{1},\mathbf{k}_{2})$ and the plane defined by the outgoing momenta $(\mathbf{k_{3}},\mathbf{k_{4}})$.
Figure 2 shows the dimensionless spin-symmetric and antisymmetric Landau parameters $F_{l}^{s,a}$ ($l=0,1,2$) as functions of $r_{s}$ . These parameters, which characterize fundamental many-body interactions in the electron gas, are defined as Legendre components of the dimensionless quasiparticle interaction $F^{s,a}(\theta)=(F_{\uparrow\uparrow}(\theta)\pm F_{\uparrow\downarrow}(\theta))/2$ with $F_{\sigma\sigma^{\prime}}(\theta)=N_{F}^{*}f_{\sigma\sigma^{\prime}}(\theta)$, where $N_{F}^{*}=\frac{m^{*}}{m}N_{F}$ is the renormalized density of states at the Fermi surface.
One pivotal observation is the linear decrease of the compressibility-related parameter: $F_{0}^{s}\approx-0.17r_{s}$. This behavior drives the Coulomb screening effect from “underscreening” to “overscreening”, with the crossover occurring at $F_{0}^{s}=-1$ where the proper compressibility approaches zero and the ion-ion interaction becomes effectively attractive at large distances [51, 46]. Our calculations show this crossover occurs near $r_{s}^{c}\approx 5.7$, which is close to the previous theoretical predictions in Ref. [45] of $r_{s}=5.25$  [^1] A further discussion of this crossover phenomenon will be discussed in Sec. IV in the supplemental material with our numerical results of the density response function and the ion-ion interaction.. Importantly, the higher-order Landau parameters remain significantly smaller than $F_{0}^{s}$, demonstrating the rapid convergence of the Legendre expansion. For comparison, Fig. 2 also shows $F_{0}^{s}$ obtained from the RPA vertex and the dimensionless zero–momentum static XC kernel $f_{\text{XC}}^{\pm}(q=0)N_{F}$ derived from the PW92 energy functional [48]. The PW92 XC kernels agree well with our numerically extracted $F_{0}^{s,a}$, confirming the approximate equality between the static zero-transfer-momentum XC kernel and the corresponding Landau parameters as
$f_{\rm XC}^{\pm}(q=0)N_{F}\approx F_{0}^{s,a},$~ (3)
for $r_{s}$ in the metallic regime—an relation specific to the homogeneous electron liquid, and not generally shared by arbitrary Fermi liquids. The theoretical insights of this relation will be discussed in Sec. III in the supplemental material.
Besides, the parameter $F_{1}^{s}$ plays a crucial role in determining the effective mass $m^{*}$ through the relation $m^{*}/m=1+F_{1}^{s}$, where $m$ represents the bare electron mass. Our finding that $F_{1}^{s}$ approaches zero in the metallic density regime, which confirms the findings through the variational Monte Carlo and VDMC [53, 54, 31] that the effective mass approximately equals the bare mass in this regime.
Figure 3: Angle-resolved Landau quasiparticle interaction for $r_{s}=1,5$ with parallel and antiparallel spin. The dotted-dashed line represents the bare exchange Coulomb interaction. The larger deviation from the bare interaction for $r_{s}=5$ indicates the Coulomb screening increases as the electron density decreases.
Figure 3 displays the angle-resolved Landau quasiparticle interaction $F_{\sigma\sigma^{\prime}}(\theta)$ versus $\cos\theta$ at $r_{s}=1$ and $5$ for both parallel ($\uparrow\uparrow$) and anti-parallel spin ($\uparrow\downarrow$) configurations. The bare exchange Coulomb interaction $W_{\rm Ex}(\mathbf{k}_{1},\mathbf{k}_{2})=-N_{F}v(|\mathbf{k}_{1}-\mathbf{k}_{2}|)$ is shown for comparison. In the parallel spin channel, many-body screening substantially modifies the quasiparticle interaction from the bare exchange. At $r_{s}=1$, $F_{\uparrow\uparrow}(\theta)$ deviates from the bare exchange primarily at small angles with rapid variation as $\theta\to\pi$, while at $r_{s}=5$, deviations occur across all angles with a smoother angular dependence. This evolution indicates increasing locality of the quasiparticle interaction and enhanced screening at larger $r_{s}$ values. For the anti-parallel spin channel, where the bare two-electron short-range interaction vanishes, the quasiparticle interaction emerges entirely from many-body effects. Notably, $F_{\uparrow\downarrow}(\theta)$ exhibits minima at $\theta=\pi$ for both $r_{s}$ values, where the incoming electron’s momentum and spin are opposite, potentially relating to Cooper pairing instability from dynamic screening potentieffects [55].
Figure 4: The dimensionless two-electron scattering amplitude $A^{\sigma\sigma^{\prime}}(\theta,\phi)$ for parallel ($\uparrow\uparrow$, top row) and anti-parallel ($\uparrow\downarrow$, bottom row) spins at $r_{s}=1$ and $r_{s}=5$ at $T/T_{F}=0.025$. Panels (a-d) show the forward scattering channel ($\phi=0$). Panels (e-f) show the Cooper channel ($\theta=\pi$). Our VDMC results (symbols) are compared against serval theoretical models. We also plot the sKO+ conjectured through analyzing the deviation of our VDMC results and the KO+ interaction. which shows good agreement with the numerical results. For $r_{s}=5$, numerical results at $T/T_{F}=0.05$ are also ploted. The divergence of the scattering amplitude at the two effectively low temperature in the Cooper channel is consistent with the expected logarithmic divergence of the four-point vertex function as $T\to 0$ [55].
Figure 4 presents the dimensionless two-electron scattering amplitude, $A^{\sigma\sigma^{\prime}}(\theta,\phi)=N_{F}^{*}a_{\sigma\sigma^{\prime}}(\theta,\phi)$, for the uniform electron gas. We analyse the forward-scattering ($\phi=0$) and Cooper ($\theta=\pi$) channels at $r_{s}=1$ and 5 for $T/T_{F}=0.025$, and include $T/T_{F}=0.05$ for $r_{s}=5$. Comparison of these datasets demonstrates that our results have reached the zero-temperature limit within error bars, except for the Cooper channel divergence caused by the Tolmachev sigularity [55]. Crucially, as this singularity is confined to a phase space of measure zero, it has a negligible impact on Fermi-surface integrals in general. Based on our numerical results, we conjecture an effective interaction (sKO+), which is more accurate than the common effective interaction, for example, the RPA and KO. This ansatz is based on the charge KO interaction within the LDA, $f^{+}_{\rm XC}(q)\approx f_{\rm XC}^{+}(0)$($f_{\rm XC}^{+}$ below), supplemented by an s-wave correction $\delta R$ for the antiparallel spin configuration:
$R^{\sigma\sigma^{\prime}}_{\mathbf{k_{1}}\mathbf{k_{2}}\mathbf{q}}\!$ $=\!\frac{v_{q}+f_{\rm XC}^{+}}{1-[v_{q}+f_{\rm XC}^{+}(0)]\Pi_{0}(q)}\!-\!f_{\rm XC}^{+}\!+\!\delta R^{\sigma\sigma^{\prime}}$ (4)
$\delta R^{\sigma\sigma^{\prime}}_{\mathbf{k_{1}}\mathbf{k_{2}}\mathbf{q}}\!\!$ $=\!(\!1\!-\!\delta_{\sigma\sigma^{\prime}}\!)\!\left[C_{0}\!+\!C_{2}\frac{(\mathbf{k}\!_{1}\!-\!\mathbf{k}\!_{2})^{2}\!+\!(\mathbf{k}\!_{1}\!-\!\mathbf{k}\!_{2}\!-\!2\mathbf{q})^{2}}{2}\right].$ (5)
The form of this correction term $\delta R$ is derived from a detailed residual analysis. We performed a polynomial expansion on the residual difference between our numerical data and the KO+ ansatz (see Sec V in the supplemental Material), which indicates that such corrections are necessary only for the antiparallel spin channel. Consequently, we formulate $\delta R$ using an ansatz inspired by effective field theory for dilute Fermi gases [56, 57], identifying the expansion coefficients $C_{0}$ and $C_{2}$ as the s-wave scattering length and effective range. With this correction, the sKO+ ansatz effectively describes the full electron–electron scattering amplitude across all tested densities and channels as shown in Fig. 4.
To highlight the advantages of the proposed sKO+ ansatz, we also benchmark common theoretical models against our high-precision VDMC data in Fig. 4, revealing the significant inadequacies of standard approximations. (i) RPA: The simplest case of Eq. (4) where $f_{\rm XC}^{+}(0)=0$ and $\delta R=0$—is clearly inadequate, especially in the low-density regime ($r_{s}\sim 5$). (ii) KO+: The charge-channel KO interaction, which corresponds to $\delta R=0$ in Eq. (4), agrees well with numerical results in the forward-scattering channel but deviates significantly approaching the Cooper channel. (iii) KO±: The full KO interaction, which further incorporates the spin channel interaction $R_{-}=\hat{\sigma}\cdot\hat{\sigma}^{\prime}\left[\frac{f_{\rm XC}^{-}}{1-f_{\rm XC}^{-}\Pi_{0}(q)}-f_{\rm XC}^{-}\right]$, performs worse as $r_{s}$ increases, particularly for antiparallel spins (Fig. 4(d)(f)). The superior performance of KO+ over KO± suggests that spin interactions are implicitly captured via the exchange component of the charge interaction, without requiring explicit spin-channel terms. (iv) The s-p approximation [18, 19]: It truncates scattering to s- and p-wave channels, which matches our numerical data well, confirming that higher-order partial waves are negligible. However, this approximation is restricted to the Fermi surface and fails to provide a complete effective interaction in other energy scales. In summary, the sKO+ ansatz resolves the deficiencies of these conventional models, offering a robust effective interaction that is formally applicable beyond the Fermi surface.
We now apply our calculated scattering amplitudes to compute macroscopic transport properties [6, 58, 10], enabling a direct validation of our numerical results against experimental benchmarks. In simple metals, deviations from the Wiedemann–Franz law—which asserts that the electrical resistivity $\rho$ and thermal resistivity $W$ relate to a universal constant $L_{0}$ via the equation $\rho/(WT)=L_{0}$—are experimentally observed, particularly above the Debye temperature [59]. This deviation manifests as an additional thermal resistance with a linear temperature dependence, identified as the electron–electron scattering contribution, $W_{ee}$ [21, 22, 23, 24]. However, a precise theoretical description has remained elusive: conventional first-principles DFT typically struggles to capture such dynamical scattering processes, while standard many-body approximations like RPA and the original KO± interaction fail to reproduce the experimental magnitude—yielding significant underestimates and overestimates, respectively (Fig. 1). By contrast, utilizing the full angular dependence of our ab initio scattering amplitudes  [^2] See Sec. SVI in the supplemental material, our VDMC calculations achieve excellent agreement with measurements for simple metals (Al, Na, K, Rb) , where band mass and core polarization corrections are negligible. Remarkably, this agreement is faithfully reproduced by the computationally efficient sKO+ ansatz, highlighting its promise for incorporating accurate e-e scattering into first-principles transport calculations of materials. Complementing these macroscopic resistivity results, we also present the quasiparticle scattering rate in the inset of Fig. 1, providing the fundamental timescale governing these transport dynamics [6, 10].
Disccusion.—In this work, we present high-precision ab initio calculations of the four-point vertex function for the 3D uniform electron gas using our developed VDMC framework. We extract Landau parameters that validate the local density approximation and reveal how the dominant $F_{0}^{s}$ parameter drives the system from underscreening toward overscreening as electron density decreases. Our analysis of electron-electron scattering amplitudes on the Fermi surface enables us to establish a simple yet accurate effective model, the sKO+ ansatz, which is based on the charge KO interaction in LDA with a small s-wave correction. The sufficiency of such short-range corrections parallels recent findings that imperfect screening at short distances necessitates such corrections to capture the physics of the low-density regime. [61]. The excellent agreement between experimental thermal resistivity measurements, our numerical data, and this effective model demonstrates that our results provide a robust foundation for first-principles transport calculations for the electron liquids in simple metals. Looking forward, our methods offer crucial numerical benchmarks for emerging experimental techniques that probe two-particle correlations in materials [15, 16, 17]. From the theoretical point of view, the sKO+ effective interaction will inspire us to construct a better low-energy effective field theory for the fermionic system, consistent with recent insights into the critical role of short-range interactions in low-density metals.
Acknowledgement—K. C. was supported by the National Key Research and Development Program of China, Grant No. 2024YFA1408604, and the National Natural Science Foundation of China under Grants No. 12047503. Z.L, P.H., and Y.D. were supported by the National Natural Science Foundation of China (under Grant No. 12275263), the Innovation Program for Quantum Science and Technology (under grant No. 2021ZD0301900), the Natural Science Foundation of Fujian Province of China (under Grant No. 2023J02032).